# The KSBA moduli space of stable log Calabi–Yau surfaces

### v2

@@ 2.2. Symington polytopes
Associated to a polarized log Calabi–Yau surface $(Y,D,L)$, there is a so called Symington “polytope” $P$, given by an integral affine manifold with singularities, which corresponds to the base of an almost toric fibration $Y\to P$ [91, §$2.2$]. Engel and Friedman showed in [31] that one can make nice choices of Symington polytopes, so that they are obtained from a toric momentum polytope by cutting out non-overlapping triangles with one edge adjacent to the boundary. In this section we review this construction, following [31, §$5.1$].
@@
We describe the construction of a Symington polytope for a polarized log Calabi–Yau surface $(Y,D,L)$ endowed with a good toric model $(\overline{Y},\overline{D},\overline{L})$ as in Definition 2.7. We denote by $\overline{P}$ the toric momentum polytope associated to the ample line bundle $\overline{L}$ on the toric surface $(\overline{Y},\overline{D})$. The edges $(\overline{P}_{i})_{1\leq i\leq n}$ of the boundary $\partial\overline{P}$ of $\overline{P}$ are naturally in one-to-one correspondence with the toric irreducible components $(\overline{D}_{i})_{1\leq i\leq n}$ of $\overline{D}$. Moreover, for every $1\leq i\leq n$, the integral length $\ell(\overline{P}_{i})$ of the edge $\overline{P}_{i}$ is given by the intersection number $\overline{L}\cdot\overline{D}_{i}$.
@@
By the proof of [31, Proposition 5.8], one can find an integral point $p_{ij}$ on $\overline{L}_{ij}$. Let $\Delta_{ij}$ be the triangle with base $\overline{P}_{ij}$ and third vertex $p_{ij}$. By construction, $\Delta_{ij}$ is an integral triangle, with base $\overline{P}_{ij}$ equal to $m_{ij}$, and with integral height also equal to $m_{ij}$. We refer to $m_{ij}$ as the size of $\Delta_{ij}$. We then define the Symington polytope $P$ associated to $(Y,D,L)$ as the integral affine manifold with singularities $P$ obtained from complement $\overline{P}\setminus\cup_{ij}\mathrm{Int}(\Delta_{ij})$, by gluing the edges adjacent to $p_{ij}$ using a matrix conjugate to
@@
This amounts to introducing a particular integral affine singularity, known as a focus-focus singularity, at each point $p_{ij}$. The corresponding monodromy invariant direction is the line $L_{ij}$ parallel to the edge $\overline{P}_{i}$ and passing through $p_{ij}$. The discriminant locus of the affine structure is given by the union of points
@@
Figure 2.1. A Symington polytope for a polarized log Calabi–Yau surface $(Y,D,L)$ as in Construction 2.10.
We refer to the vertex $p_{ij}$ of $\Delta_{ij}$ as the interior vertex of $\Delta_{ij}$, and to the edges of $\Delta_{ij}$ adjacent to $p_{ij}$ as the interior edges of $\Delta_{ij}$. Note that, sometimes the interior vertices of several triangles $\Delta_{ij}$ may coincide, leading to more complex singularities on $Y$ than focus-focus singularities.
The following result shows that the monodromy invariant lines have always a non-trivial intersection with the Symington polytope $P$.
Lemma 2.11.
For every $1\leq i\leq n$, $1\leq j\leq N_{i}$, the intersection of the monodromy invariant line $L_{ij}$ with the interior of $P$ is an interval of positive length.
Proof.
Denote by $\ell_{ij}=0$ the affine equation of $\overline{P}_{i}$. Then $L_{ij}$ is the line of equation $\ell_{ij}=m_{ij}$. Denote by $d_{ij}$ the greatest integral distance between the line $L_{ij}$ and a point $p\in P\cap\{\ell_{ij}\geq m_{ij}\}$. To prove Lemma 2.11, it is enough to show that $d_{ij}>0$.
We will prove that $d_{ij}>0$ by giving a geometric interpretation of $d_{ij}$. Let $(\overline{Y}^{\prime},\overline{D}^{\prime})$ be the toric variety obtained from $(\overline{Y},\overline{D})$ by adding to its fan the ray opposite to the ray corresponding to $\overline{P}_{i}$. These two opposite rays in the fan define a generically $\mathbb{P}^{1}$-fibration $\overline{Y}^{\prime}\rightarrow\mathbb{P}^{1}$. Denote by $(\tilde{Y}^{\prime},\tilde{D}^{\prime})$ the corresponding blow-up of $(Y,D)$, with a generically $\mathbb{P}^{1}$-fibration $\tilde{Y}^{\prime}\rightarrow\mathbb{P}^{1}$. Then, the pre-image of $E_{ij}$ is an irreducible component $E_{ij}^{\prime}$ of a fiber of this map. Denote by $F_{ij}$ the other irreducible component of this fiber and $g^{\prime}:\widetilde{Y}^{\prime}\rightarrow Y$ the contraction induced by $g$. Then, we have $d_{ij}=L\cdot g^{\prime}_{\star}F_{ij}$. Since $g^{\prime}$ is a corner blow-up, we have $g^{\prime}_{\star}F_{ij}\neq 0$, and so we conclude that $d_{ij}=L\cdot g^{\prime}_{\star}F_{ij}>0$ since $L$ is ample. ∎
@@ 2.3. Good polyhedral decompositions
Let $(Y,D,L)$ be a polarized log Calabi–Yau surface with associated Symington polytope $P$, obtained from a toric polytope $\overline{P}$ as in Construction 2.10. In this section, we first define in Definition 2.13 the notion of a good polyhedral decomposition on $\overline{P}$, which is a particular regular polyhedral decomposition appropriately compatible with the cuts producing the Symington polytope. Then, we prove the existence of such good polyhedral decompositions in Theorem 2.14.
@@
We denote by $\overline{\mathscr{P}}^{[k]}$ the set of k-dimensional cells of $\overline{\mathscr{P}}$. A polyhedral decomposition $\overline{\mathscr{P}}$ on $\overline{P}$ is called regular if there exists a convex piecewise linear function $\varphi:\overline{P}\rightarrow\mathbb{R}$ such that the polyhedra $\{\sigma\}$ of $\overline{\mathscr{P}}$ corresponds to the domain of linearity of $\varphi$. In this situation, we refer to $\overline{\mathscr{P}}$ as the decomposition induced by $\varphi$. In what follows, we describe good regular polyhedral decompositions on $\overline{P}$. For this we first need the following observation.
@@
Without loss of generality, assume coordinates for the vertex $p_{ij}$ are $(0,0)$, and for the other two vertices of $\Delta_{ij}$ on $\overline{P}_{ij}$ are $(a,b)$ and $(a,c)$ respectively, with $b>c$. Since the edge of $\Delta_{ij}$ on the side $\overline{P}_{ij}$ has the same integral length as the height of $\Delta_{ij}$ by Construction 2.10, we have $a=b-c$. Then, we obtain $d:=gcd(a,b)=gcd(a,c)$, and the sets of integral points on the edges connecting $p_{ij}=(0,0)$ to $(a,b)$ and to $(a,c)$ respectively have the same cardinalities, equal to $d+1$:
[…]
Now we call the decomposition of $\Delta_{ij}$, into the union of the triangle $\Delta_{ij,0}$ whose set of vertices is
@@
and the $(d-1)$ trapezoids $\Delta_{ij,k}$ whose set of vertices are
@@
First, start with the polyhedral subdivision $\overline{\mathscr{P}}_{0}$ of $\overline{P}$ given by the union of the triangles $\overline{T}_{i}$. Let $\overline{\mathscr{P}}_{1}$ be the polyhedral decomposition obtained by adding to $\overline{\mathscr{P}}_{0}$ the edges joining $p$ to all vertices of the edges of $\Delta_{ij}$’s – this is possible to do without any intersection with the edges of $\overline{\mathscr{P}}_{0}$ since Construction 2.10 ensures that the triangles $\Delta_{ij}$’s lie inside triangles $\overline{T}_{ij}$’s which intersect at most along common edges. Note that the polyhedral decomposition $\overline{\mathscr{P}}_{1}$ of $\overline{P}$ is regular – see Figure 2.1, and so there exists a convex PL function $\varphi_{1}:\overline{P}\to\mathbb{R}$ inducing $\overline{\mathscr{P}}_{1}$.
@@
Recall that, by Proposition 2.12, there are the same number integral points on each of the two interior edges of $\Delta_{ij}$’s. Let $\{a^{1}_{ij},\ldots,a^{d}_{ij}\}$ and $\{b^{1}_{ij},\ldots,b^{d}_{ij}\}$ respectively denote the set of integral points except from the interior vertex $p_{ij}$, on the interior edges of $\Delta_{ij}$. Let ${\epsilon_{ij}^{1},\ldots,\epsilon_{ij}^{d}}$ be a set of positive real numbers such that the points
$(p_{ij},\varphi_{0}(p_{ij}))\,,(a_{ij}^{1},\varphi_{0}(a_{ij}^{1})-\epsilon_{ij}^{1})\,,...,\,,(a_{ij}^{d},\varphi_{0}(a_{ij}^{d})-\epsilon_{ij}^{d})$ (2.6)
@@
are exactly the bending points of a PL function on the interior edge of $\Delta_{ij}$ containing the points $a_{ij}^{k}$’s. As $\varphi_{1}$ is affine on $\Delta_{ij}$, it follows that the points
$(p_{ij},\varphi_{0}(p_{ij}))\,,(b_{ij}^{1},\varphi_{0}(b_{ij}^{1})-\epsilon_{ij}^{1})\,,...,\,,(b_{ij}^{d},\varphi_{0}(b_{ij}^{d})-\epsilon_{ij}^{d})$ (2.7)
@@
are also exactly the bending points of a PL function on the interior edge of $\Delta_{ij}$ containing the points $b_{ij}^{k}$’s. Let $\varphi_{2}:\overline{P}\rightarrow\mathbb{R}$ be the convex PL function whose graph is the lower convex hull of the points in (2.6), the points in (2.7), and the points $(v,\varphi(v))$, where $v$ are vertices of $\overline{\mathscr{P}}_{1}$ not contained in triangles $\Delta_{ij}$. Denote by $\overline{\mathscr{P}}_{2}$ the regular decomposition of $\overline{P}$ induced by $\varphi_{2}$. If $\epsilon_{ij}^{1},\cdots,\epsilon_{ij}^{d}$ are small enough for all $1\leq i\leq n$, $1\leq j\leq N_{i}$, then $\overline{\mathscr{P}}_{2}$ satisfies condition (i) of Definition 2.13.
@@
To obtain a good decomposition, that also satisfies condition (ii) of Definition 2.13, we start with $\overline{\mathscr{P}}_{2}$ and then successively pull generically all the integral points $v$ of $\overline{P}$ which are not vertices of $\overline{\mathscr{P}}_{2}$, that is we successively modify the convex function by lowering its value at $v$ by a small general enough positive amount – see for instance [27, §4.3]. The resulting decomposition $\overline{\mathscr{P}}$, illustrated in Figure 2.2, coincides with $\overline{\mathscr{P}}_{2}$ on each triangle $\Delta_{ij}$, and so also satisfies condition (i) of Definition 2.13. Moreover, on the complement in $\overline{P}$ of the interior of the triangles $\Delta_{ij}$, $\overline{\mathscr{P}}$ is a triangulation containing all integral points, and so $\overline{\mathscr{P}}$ satisfies condition (ii) of Definition 2.13, since by Pick’s formula an integral triangle containing no integral point apart from its vertices is necessarily a triangle of size 1.

### v3

@@ 2.2. Symington polytopes
Given a generic polarized log Calabi–Yau surface $(Y,D,L)$, one can associate a Symington “polytope” $P$, which is an integral affine manifold with singularities. This polytope serves as the base of an almost toric fibration $Y\to P$ [95, §$2.2$]. Engel–Friedman demonstrated in [33] that well-chosen Symington polytopes can be obtained from a toric momentum polytope by cutting out non-overlapping triangles. In this section we review this construction, following [33, §$5.1$].
@@
We describe the construction of a Symington polytope for a generic polarized log Calabi–Yau surface $(Y,D,L)$ endowed with a good toric model $(\overline{Y},\overline{D},\overline{L})$ as in Definition 2.7. We denote by $\overline{P}$ the toric momentum polytope associated to the ample line bundle $\overline{L}$ on the toric surface $(\overline{Y},\overline{D})$. The edges $(\overline{P}_{i})_{1\leq i\leq n}$ of the boundary $\partial\overline{P}$ of $\overline{P}$ are naturally in one-to-one correspondence with the toric irreducible components $(\overline{D}_{i})_{1\leq i\leq n}$ of $\overline{D}$. Moreover, for every $1\leq i\leq n$, the integral length $\ell(\overline{P}_{i})$ of the edge $\overline{P}_{i}$ is given by the intersection number $\overline{L}\cdot\overline{D}_{i}$.
@@
By the proof of [33, Proposition 5.8], there exists an integral point $p_{ij}$ on $\overline{L}_{ij}$. Let $\Delta_{ij}$ be the triangle with base $\overline{P}_{ij}$ and third vertex $p_{ij}$. By construction, $\Delta_{ij}$ is an integral triangle, with its base $\overline{P}_{ij}$ of integral length $m_{ij}$, and integral height also equal to $m_{ij}$. We refer to $m_{ij}$ as the size of $\Delta_{ij}$. We then define the Symington polytope $P$ associated to $(Y,D,L)$ as the integral affine manifold with singularities $P$ obtained from complement $\overline{P}\setminus\cup_{ij}\mathrm{Int}(\Delta_{ij})$, by gluing the edges adjacent to $p_{ij}$ using a matrix conjugate to
@@
This process introduces a specific type of integral affine singularity, known as a focus-focus singularity, at each point $p_{ij}$. The corresponding monodromy invariant direction is the line $L_{ij}$ which is parallel to the edge $\overline{P}_{i}$ and passes through $p_{ij}$. The discriminant locus of the affine structure is given by the union of points
@@
Figure 1. A Symington polytope for a polarized log Calabi–Yau surface $(Y,D,L)$ as in Construction 2.10
We refer to the vertex $p_{ij}$ of $\Delta_{ij}$ as the interior vertex of $\Delta_{ij}$, and to the edges of $\Delta_{ij}$ adjacent to $p_{ij}$ as the interior edges of $\Delta_{ij}$ – see Figure 1. Note that in some cases, the interior vertices of several triangles $\Delta_{ij}$ may coincide, resulting in more complex integral affine singularities on $Y$ beyond focus-focus singularities. The following result establishes that the area of the Symington polytope $P$ is given by $\frac{1}{2}L^{2}$, a positive quantity due to the ampleness of the line bundle $L$. Consequently, $P$ is non-empty and two-dimensional.
Lemma 2.11.
Let $P$ be a Symington polytope associated to a polarized log Calabi–Yau surface $(Y,D,L)$ as in Construction 2.10. Then, the area of $P$ is equal to $\frac{1}{2}L^{2}$.
Proof.
We use the notation introduced in Construction 2.10. First, since $g^{\star}L=\overline{L}-\sum_{i,j}m_{ij}E_{ij}$, it follows that $L^{2}=\overline{L}^{2}-\sum_{i,j}m_{ij}^{2}$. On the other hand, since $P$ is obtained from $\overline{P}$ by removing the non-overlapping triangles $\Delta_{ij}$ of sizes $m_{ij}$, we obtain $\mathrm{Area}(P)=\mathrm{Area}(\overline{P})-\frac{1}{2}\sum_{i,j}m_{i,j}^{2}$. Finally, by standard toric geometry [39, §5.3, p111, Corollary], we have $\mathrm{Area}(\overline{P})=\frac{1}{2}\overline{L}^{2}$, and so the result follows. ∎
Figure 2. Illustration of the first part of the proof of Lemma 2.12
The following result shows that the monodromy invariant lines have always a non-trivial intersection with the interior of $\overline{P}$.
Lemma 2.12.
Let $P$ be a Symington polytope associated to a polarized log Calabi–Yau surface $(Y,D,L)$ as in Construction 2.10. Then, for every $1\leq i\leq n$, $1\leq j\leq N_{i}$, the intersection of the monodromy invariant line $L_{ij}$ with the interior of $P$ is an interval of positive length.
Proof.
We first prove the result when $\Delta_{ij}$ is the only cut triangle. Let $\ell_{i}=0$ be the affine equation defining $\overline{P}_{i}$, so that $L_{ij}$ is the line of equation $\ell_{i}=m_{ij}$. Define $d_{ij}$ the greatest integral distance between the line $L_{ij}$ and a point $q\in\overline{P}\cap\{\ell_{i}\geq m_{ij}\}$. To prove Lemma 2.12, it suffices to show that $d_{ij}>0$. Indeed, in this situation, let $q$ be a point in $\overline{P}$ with $\ell_{i}(q)>m_{ij}$, and denote by $p_{1}$ and $p_{2}$ the two endpoints of the line segments $\overline{P}_{ij}$. Since the points $p_{1},p_{2},p_{ij}$ are not collinear, at least one of the triangles with vertices $p_{1},p_{ij},q$ or $p_{2},p_{ij},q$ must be non-flat – see Figure 2. Such a non-flat triangle necessarily has a non-trivial intersection with $L_{ij}$ and is contained within $\overline{P}$, as the polygon $\overline{P}$ is convex and contains the points $p_{1}$, $p_{2}$, $p_{ij},q$.
@@
Finally, we prove Lemma 2.12 in the general case where $\Delta_{ij}$ is not the only cut triangle. If $\Delta_{ij}$ has an interior edge that is not entirely contained within another cut triangle, then by the first part of the proof, the corresponding segment of $L_{ij}$ has a non-trivial intersection with $P$. From Construction 2.10, the only way both interior edges of $\Delta_{ij}$ are fully contained within other cut triangles is when $p_{ij}$ coincides with the central point $p$. In this case, one can attempt to propagate $L_{ij}$ through the neighboring cut triangles to obtain a non-trivial intersection with $P$. This will always work unless the cut triangles completely cover $\overline{P}$. However, this does not occur, as the Symington polytope $P$ has positive area by Lemma 2.11 and $L^{2}>0$ due to the ampleness of $L$. This concludes the proof of Lemma 2.12. ∎
@@ 2.3. Good polyhedral decompositions
Let $(Y,D,L)$ be a generic polarized log Calabi–Yau surface with associated Symington polytope $P$, obtained from a toric polytope $\overline{P}$ as in Construction 2.10. In this section, we first define in Definition 2.15 the notion of a good polyhedral decomposition on $\overline{P}$, which is a particular regular polyhedral decomposition appropriately compatible with the cuts producing the Symington polytope. Then, we prove the existence of such good polyhedral decompositions in Theorem 2.16.
@@
We denote by $\overline{\mathscr{P}}^{[k]}$ the set of k-dimensional cells of $\overline{\mathscr{P}}$. A polyhedral decomposition $\overline{\mathscr{P}}$ on $\overline{P}$ is called regular if there exists a convex piecewise linear function $\varphi:\overline{P}\rightarrow\mathbb{R}$ such that the polyhedra $\{\sigma\}$ of $\overline{\mathscr{P}}$ corresponds to the domain of linearity of $\varphi$. In this situation, we refer to $\overline{\mathscr{P}}$ as the decomposition induced by $\varphi$. In what follows, we describe good regular polyhedral decompositions on $\overline{P}$. We first need the following observation.
@@
Without loss of generality, we can assume that the vertex $p_{ij}$ has coordinates $(0,0)$, and that the other two vertices of $\Delta_{ij}$ on $\overline{P}_{ij}$ have coordinates $(a,b)$ and $(a,c)$ respectively, with $b>c$. Since the edge of $\Delta_{ij}$ on the side $\overline{P}_{ij}$ has the same integral length as the height of $\Delta_{ij}$ by Construction 2.10, we have a=bca=b-c. Then, we obtain $d:=\gcd(a,b)=\gcd(a,c)$, and the sets of integral points on the edges connecting $p_{ij}=(0,0)$ to $(a,b)$ and to $(a,c)$ respectively have the same cardinality, equal to $d+1$:
$\left|\left\{\left(\frac{ka}{d},\frac{kb}{d}\right)\,\ |\,\ 0\leq k\leq d\right\}\right|=\left|\left\{\left(\frac{ka}{d},\frac{kc}{d}\right)\,\ |\,\ 0\leq k\leq d\right\}\right|=d+1.$
Definition 2.14.
The standard decomposition of $\Delta_{ij}$ is the polyhedral decomposition of $\Delta_{ij}$ consisting of:
  • i)
    the triangle $\Delta_{ij,0}$ whose set of vertices is
@@
  • ii)
    the (d1)(d-1) trapezoids $\Delta_{ij,k}$ whose set of vertices are
@@
We will construct regular polyhedral decompositions of $\overline{P}$ by successively pulling of integral vertices. This process involves iteratively modifying a piecewise linear convex function by lowering its value at integral vertices by a small enough positive amount – see for instance [28, §4.3].
Let $\overline{\mathscr{P}}_{1}$ be the regular polyhedral decomposition of $\overline{P}$ obtained from the trivial decomposition by first pulling the central point $p$, and then pulling the vertices of the triangles $\Delta_{ij}$ – see Figure 3. Denote by $\varphi_{1}$ a convex piecewise linear function $\varphi_{1}:\overline{P}\to\mathbb{R}$ inducing $\overline{\mathscr{P}}_{1}$.
@@
By Proposition 2.13, there are the same number of integral points on each of the two interior edges of $\Delta_{ij}$’s. Let $\{a^{1}_{ij},\ldots,a^{d}_{ij}\}$ and $\{b^{1}_{ij},\ldots,b^{d}_{ij}\}$ respectively denote the set of integral points on the interior edges of $\Delta_{ij}$, except for the interior vertex $p_{ij}$. Let $\epsilon_{ij}^{1},\ldots,\epsilon_{ij}^{d}$} be a set of positive real numbers such that the points
$(p_{ij},\varphi_{0}(p_{ij}))\,,(a_{ij}^{1},\varphi_{0}(a_{ij}^{1})-\epsilon_{ij}^{1})\,,\cdots,\,,(a_{ij}^{d},\varphi_{0}(a_{ij}^{d})-\epsilon_{ij}^{d})$ (2.6)
@@
are exactly the bending points of a piecewise linear function on the interior edge of $\Delta_{ij}$ containing the points $a_{ij}^{k}$’s. Since $\varphi_{1}$ is affine on $\Delta_{ij}$, the points
$(p_{ij},\varphi_{0}(p_{ij}))\,,(b_{ij}^{1},\varphi_{0}(b_{ij}^{1})-\epsilon_{ij}^{1})\,,\cdots,\,,(b_{ij}^{d},\varphi_{0}(b_{ij}^{d})-\epsilon_{ij}^{d})$ (2.7)
@@
are also exactly the bending points of a piecewise linear function on the interior edge of $\Delta_{ij}$ containing the points $b_{ij}^{k}$’s. Let $\varphi_{2}:\overline{P}\rightarrow\mathbb{R}$ be the convex piecewise linear function whose graph is the lower convex hull of the points in (2.6), the points in (2.7), and the points $(v,\varphi(v))$, where $v$ are vertices of $\overline{\mathscr{P}}_{1}$ not contained in triangles $\Delta_{ij}$. Denote by $\overline{\mathscr{P}}_{2}$ the regular decomposition of $\overline{P}$ induced by $\varphi_{2}$. If $\epsilon_{ij}^{1},\cdots,\epsilon_{ij}^{d}$ are small enough for all $1\leq i\leq n$, $1\leq j\leq N_{i}$, then $\overline{\mathscr{P}}_{2}$ satisfies condition i) of Definition 2.15.
@@
To obtain a good decomposition that also satisfies condition ii) of Definition 2.15, we start with $\overline{\mathscr{P}}_{2}$ and then successively pull generically all the integral points $v$ of $\overline{P}$ which are not vertices of $\overline{\mathscr{P}}_{2}$. The resulting decomposition $\overline{\mathscr{P}}$, illustrated in Figure 3, coincides with $\overline{\mathscr{P}}_{2}$ on each triangle $\Delta_{ij}$, and so also satisfies condition i) of Definition 2.15. Moreover, on the complement in $\overline{P}$ of the interior of the triangles $\Delta_{ij}$, $\overline{\mathscr{P}}$ is a triangulation containing all integral points, and so $\overline{\mathscr{P}}$ satisfies condition ii) of Definition 2.15, since by Pick’s formula an integral triangle containing no integral point apart from its vertices is necessarily a triangle of size 1.
Figure 3. The left hand figure illustrates $\overline{\mathscr{P}}_{2}$. To obtain the final decomposition $\overline{\mathscr{P}}$ the red edges are inserted, as in the right hand figure.
2.4. Maximal degenerations
Given a generic polarized log Calabi–Yau surface $(Y,D,L)$ with an associated Symington polytope $P$, obtained from $\overline{P}$ as described in Construction 2.10, we naturally obtain a polyhedral decomposition $\mathscr{P}$ on $P$ by restricting the good polyhedral decomposition $\overline{\mathscr{P}}$ on $\overline{P}$ to $P$. In this section, we outline the construction of a maximal degeneration of $(Y,D,L)$, meaning that the central fiber contains a 0-dimensional stratum. Here, stratum refers to the intersection of irreducible components of the central fiber.
Let $\varphi$ be a convex piecewise linear function on $\overline{P}$ defining the regular polyhedral decomposition $\overline{\mathscr{P}}$. We will construct a one-parameter maximal polarized degeneration $(\mathcal{Y}_{\varphi},\mathcal{D}_{\varphi},\mathcal{L}_{\varphi})\rightarrow\mathbb{A}^{1}$, with general fiber deformation equivalent to $(Y,D,L)$.
We begin with the Mumford toric degeneration $(\overline{\mathcal{Y}}_{\varphi},\overline{\mathcal{D}}_{\varphi},\overline{\mathcal{L}}_{\varphi})\rightarrow\mathbb{A}^{1}$ of the polarized toric surface $(\overline{Y},\overline{D},\overline{L})$ induced by the polyhedral decomposition $\overline{\mathscr{P}}$ of $\overline{P}$ [42, Example 3.6]. The intersection complex of the central fiber $(\overline{Y}_{\varphi,0},\overline{D}_{\varphi,0},\overline{L}_{\varphi,0})$ is the polyhedral decomposition $\overline{\mathscr{P}}$ of $\overline{P}$. The irreducible components of the central fiber are the polarized toric surfaces $Y^{f}$ with momentum polytopes the 2-dimensional faces $f$ of $\overline{\mathscr{P}}$ and they are glued together along the toric divisors $Y^{e}$ corresponding to the edges $e$ of $\overline{\mathscr{P}}$.
Let $h_{\varphi}:\widetilde{\mathcal{Y}}_{\varphi}\rightarrow\overline{\mathcal{Y}}_{\varphi}$ be the blow-up of $\overline{\mathcal{Y}}_{\varphi}$ along the sections $s_{ij}$. We denote by $\widetilde{\mathcal{D}}_{\varphi}$ the strict transform of $\overline{\mathcal{D}}_{\varphi}$, and by $\widetilde{\mathcal{L}}_{\varphi}$ the line bundle on $\widetilde{\mathcal{Y}}_{\varphi}$ defined by
$\widetilde{\mathcal{L}}_{\varphi}=h_{\mathscr{P}}^{\star}\overline{\mathcal{L}}_{\varphi}-\sum_{i=1}^{n}\sum_{j=1}^{n}m_{ij}\mathcal{E}_{ij}\,$,